Problem 4.4

You got asked to analyze a 5-year project for your firm. The project produces an annual revenue of $28,000, but requires an annual labor and materials cost of $6,000. To initiate the project your firm must invest $18,000. The salvage value of the project is $0 at the end of the 5-year useful life. Use straight line depreciation and a 40% income tax rate to compute the after-tax cash flows (15 points) and the IRR for the ATCF of this project (5 points).
Answer #1

SL Dep per year = (18000 - 0) / 5 = 3600

Tax rate = 40%

Net cash flow = 28000 - 6000 = 22000

Taxable income = Net cash flow - Depreciation = 22000 - 3600 = 18400

Tax = Tax rate * Taxable income = 0.4*18400 = 7360

ATCF per yr = BTCF - Tax = 22000 - 7360 = 14640

Using Excel

Year Untaxed BTCF Taxed BTCF SL Dep Tax income Tax ATCF
0 -18000 -18000
1 22000 3600.00 18400.00 7360.00 14640.00
2 22000 3600.00 18400.00 7360.00 14640.00
3 22000 3600.00 18400.00 7360.00 14640.00
4 22000 3600.00 18400.00 7360.00 14640.00
5 0 22000 3600.00 18400.00 7360.00 14640.00
IRR 76.60%

IRR = 76.60%

Showing formula in Excel

Year Untaxed BTCF Taxed BTCF SL Dep Tax income Tax ATCF
0 -18000 =B2
1 =28000-6000 =(18000-0)/5 =C3-D3 =E3*0.4 =C3-F3
2 =28000-6000 =(18000-0)/5 =C4-D4 =E4*0.4 =C4-F4
3 =28000-6000 =(18000-0)/5 =C5-D5 =E5*0.4 =C5-F5
4 =28000-6000 =(18000-0)/5 =C6-D6 =E6*0.4 =C6-F6
5 0 =28000-6000 =(18000-0)/5 =C7-D7 =E7*0.4 =C7-F7
IRR =IRR(G2:G7)